OpenStudy (anonymous):

Given A=integrate from 0 to 1 e^(-t^2) dt B=integrate from 0 to 1/2 e^(-t^2) dt compute the following dobule integral: 2 * integrate from -1/2 to 1 [integrate from 0 to x e^(-y^2) dy]dx as a function of A and B. Check that exists positives integers m and n and: I= m*A-n*B+e^(-1)-e^(-1/4) I tried reversing the integral order of the double integral to be able to compute easier, but it gave me: e^(-1)-e^(-1/4) + f(x) + c And if you made in the given order you got: e^(-1)-e^(-1/4) + f(y) + c And f(y) is where A and B came from, so I don't know how to solve this and... (to be continued)

OpenStudy (anonymous):

And I can't compute integrate from 0 to x [e^-(y^2)] dy I think I'll try using the definition of double integral and trying to work with the sumatories, but if anyone can help me I'll be happy. PD: Both constants haven't to be the same, it was a mistake to use "c" in both e^(-1)-e^(-1/4) + f(x) + c e^(-1)-e^(-1/4) + f(y) + k should be better.

OpenStudy (anonymous):

Here are the ecuations in a easier way to read. \[A=\int\limits_{0}^{1}e^{-t^2}dt\] \[B=\int\limits_{0}^{1/2}e^{-t^2}dt\] \[I=2\int\limits_{-1/2}^{1}[\int\limits_{0}^{x}e^{-y^2}dy]dx\] \[I = mA -nB +e^{-1}-e^{-1/4}\]

OpenStudy (anonymous):

I had an idea. First I define: f(x)=\[f(x)=\int\limits_{0}^{x} e^{-y^2}dy\]e^(-y^2) I'm evaluating the integral of \[e^{-y^2}\] between 0 and x, and this antiderivate in 0 is a constant, so: \[f'(x)=e^{-x^2}\] Then I can use integration by parts: \[2\int\limits_{}^{}f(x)dx=2(xf(x)-\int\limits_{}^{}xf'(x)dx)\] \[2\int\limits_{}^{}xf'(x)dx=2\int\limits_{}^{}xe^{-x^2}dx=-e^{-x^2}\] You can solve the last part with a simple substitution, -x^2=u \[2\int\limits_{}^{}xe^{-x^2}dx=2\int\limits\limits\limits_{}^{}xe^{-u^2}/(-2x)du=-\int\limits\limits_{}^{}e^{u}du=-e^u=-e^{-x^2}\] Then: \[2\int\limits_{}^{}f(x)dx=2xf(x)-(-e^{-x^2})\] \[2\int\limits\limits_{-1/2}^{1}f(x)dx=2(1)f(1)+e^{-1^2}-(2(-1/2)f(-1/2)+e^{-(-1/2)^2})=\] \[2f(1)+e^{-1^2}+f(-1/2)-e^{-(1/4)}\] \[B=\int\limits_{-1/2}^{0}e^{-t^2}dt \rightarrow -B=\int\limits_{0}^{1/2}e^{-t^2}dt\] Because t is squared, so \[e^{-t^2}=e^{-(-t^2)}\] And the function is simetric. Then \[f(-1/2)=-f(1/2)\] \[2f(1)+e^{-1}-f(1/2)-e^{-(1/4)}=2A-B+e^{-1}-e^{-(1/4)}\] Then m=2 and n=1